OpenStudy (anonymous):

Find dy/dx by implicit differentiation for x^2+y^2=100. The answer should end up as -x/y

OpenStudy (tkhunny):

What's your plan for that? Hint: (d/dx) y = y' Hint: (d/dx) y^2 = 2yy'

OpenStudy (anonymous):

I missed the section lecture.... :/ so my best guess is too take the derivative of each so I started at 2x+2y=0

OpenStudy (anonymous):

Close... but technically, it is 2xx' + 2yy' = 0. Since x' = 1, then normally it isn't written... So you end up with 2x + 2yy' = 0 which can be solved easily for y'...

OpenStudy (anonymous):

Ok I understand, thank you both of you. I think I got the basics down. It makes much more sense when I see 2xx' and 2yy' written out now